OpenStudy (anonymous):

Find dy/dx by implicit differentiation and evaluate the derivative at the given point. tan(5x + y) = 5x

myininaya (myininaya):

(5+y')sec^2(5x+y)=5

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

sec^2(5x+y)(5+y')=5

OpenStudy (anonymous):

these guys are beasts!

OpenStudy (anonymous):

yeah..so imran how would you continue that

OpenStudy (anonymous):

.....

myininaya (myininaya):

I would distribute first 5sec^2(5x+y)+y'sec^2(5x+y)=5

myininaya (myininaya):

This is hard to do on the iPhone lol

OpenStudy (anonymous):

you need a medal for even trying lol

OpenStudy (anonymous):

lol sorryy

OpenStudy (anonymous):

\[y' \text{Sec}^2[5 x+y]+5 \text{Sec}^2[5 x+y]=5\] \[y' \text{Sec}^2[5 x+y]=5-5 \text{Sec}^2[5 x+y]\] \[y'=\frac{5-5 \text{Sec}^2[5 x+y]}{\text{Sec}^2[5 x+y]}\] \[\frac{5}{\text{Sec}^2[5 x+y]}-5\] \[-5+5\text{Cos}^2[5 x+y]\] \[-5+5\left(1-\text{Sin}^2[5x+y]\right)\] \[y'=-5 \text{Sin}^2[5 x+y]\]